• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name aniol

    Log entry time 21:16:41 on June 9,2010

    Entry number 325468

    keyword=target vacuum

    24 hours ago (entry 325325) the vacuum was better, about 2e-6. Today it is about 6.8e-6. 50uA beam and cooler is on, heater is off.


    FIGURE 1

    FIGURE 2